Can someone help me with this problem?
I need to find four solutions for the following equation:
When x number of units are sold, the price of each unit (in dollars) is given by p = 1x/2 + 75. I need to find the unit price when the following quantities are sold: 2, 7, 9, 11.
Thanks,
Sher

Answer by tutorcecilia(2152)

Plug-in each value for x and solve the equation.
Example:
P(x) = 1x/2 + 75
P(2)=1(2)/2 + 75
P(2)=2/2+75
P(2)=1+75
P(2)=76
.
Follow this same pattern for all of the other quantities: 7, 9, 11